OpenStudy (anonymous):

First question ever! Help Appreciated! Find the x-intercept and y-intercept of the line given by the equation: 7x + 5y = 9 Thanks!

OpenStudy (karatechopper):

x-int= make y =0 y-int= make x=0

OpenStudy (anonymous):

Transform the equation into the form, x/a+y/b=1, then a is the x-intercept and b the y-intercept.

OpenStudy (karatechopper):

x int: 7x+5(0)=9 y-int: 7(0)+5x=9

OpenStudy (karatechopper):

can you solve now?

OpenStudy (anonymous):

Thanks guys, this was a little demo question to see how this website works! Its pretty awesome!

OpenStudy (karatechopper):

o ok! Welcome to OpenStudy! If any help needed just ask me! i will be glad to help!